iuytyuioiuytyuiop:

Find the equation of the line with slope 2/5 and y-intercept (0,−4). (Write the equation in standard form.)

Angle:

Let's first write the equation in y-intercept form and then we can move things around to make it standard form y-intercept form looks like: `y = mx + b`

Angle:

do you know how to write the equation for this problem in y-intercept form?

iuytyuioiuytyuiop:

I don't

Angle:

ok, so `b` is equal to the y value of the y intercept coordinate that is going to be the `-4` from the (0, -4) that they gave us

Angle:

then `m` represents the slope and they tell us that the slope is `2/5`

Angle:

if we plug in b = -4 and m = 2/5 into the equation y = mx + b what would we get?

iuytyuioiuytyuiop:

y= 2/5x + -4

Angle:

perfect! now, we want to rearrange that equation into standard form standard form is when all of the numbers are whole numbers and where the x and y values are on the left side of the equals sign `ax + by = c` where a, b, and c are whole numbers

iuytyuioiuytyuiop:

ax+-4y=c

Angle:

mm not quite y= 2/5x + -4 we taking that equation and moving the 2/5x to the left side

iuytyuioiuytyuiop:

2/5x+=4y=c

iuytyuioiuytyuiop:

-4

Angle:

-2/5 x + y = -4 then I want you to multiply each part of that equation by `-5` to get rid of the fraction

Angle:

-2/5 times -5 1 times -5 -4 times -5

iuytyuioiuytyuiop:

2x+-5=20

Angle:

2x + -5y = 20 perfect! :)

iuytyuioiuytyuiop:

thanks